If uncertainty in momentum is twice the uncertainty in position of an electron then uncertainty in velocity is:



Grade:upto college level

1 Answers

Aditi Chauhan
10 years ago

(A, B)

Sol.    B.E. of 4th state = 13.6 z^2/n^2     =>    13.6 z^2/4^2  = 13.6    =>    z = 4

        sample is Be3+        =>    energy of electron in 3rd state

        state = 1.5 × 42 = 24 eV

    therefore 25 eV photon will cause ionization.